OpenStudy (anonymous):

The leg of a right triangle is 2 units and the hypotenuse is 4 units. What is the length, in units, of the other leg of the triangle? A)2 units B) 6 units C)Square root of 12 units D)Square root of 20 units

OpenStudy (alexalas):

Alright we use the Pythagorean Theorem \[a ^{2}+b ^{2}=c ^{2}\] where C is the hypotenuse to find a leg you change the formula to look like \[a^{2}=c ^{2}-b ^{2}\] can you fill in C and B

OpenStudy (anonymous):

A squared+ B squared= C squared so 2 squared+ ?+ 4 squared

OpenStudy (anonymous):

OHH so A i was thinking that but then i thought you had to multiply all of those numbers by its own number

OpenStudy (anonymous):

like 2 times 2 + 2 times 2= 4 times 4

OpenStudy (alexalas):

do you know the answer?

OpenStudy (anonymous):

either A or B :/

OpenStudy (alexalas):

nope

OpenStudy (anonymous):

wait it cant be B

OpenStudy (alexalas):

\[a^{2}=c^2-b^2\] fill in c and b c=4 and b=2 \[a^{2}=4^2-2^2\] that equals (you answer)

OpenStudy (anonymous):

oh so 2 i dont want to be wrong but sorry if i am im just lost

OpenStudy (alexalas):

its okay \[a ^{2}=16-4\] \[a ^{2}=12\] so what does a equal

OpenStudy (anonymous):

omg so we did have to multiply it omg thanks so C

OpenStudy (alexalas):

yes the answer is C